OpenStudy (anonymous):

Show that:

OpenStudy (anonymous):

\[\ln (1+e ^{2x})=2x+\ln(1+e ^{-2x})\]

OpenStudy (anonymous):

@Nnesha @dan815

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Do you need to solve for X?

OpenStudy (anonymous):

@Andrew12840

OpenStudy (anonymous):

no prove it

OpenStudy (michele_laino):

please note that: \[\begin{gathered} \ln \left( {1 + {e^{ - 2x}}} \right) = \ln \left( {1 + \frac{1}{{{e^{2x}}}}} \right) = \hfill \\ = \ln \left( {\frac{{1 + {e^{2x}}}}{{{e^{2x}}}}} \right) \hfill \\ \end{gathered} \]